A man flies an airplane from Fargo to Bismark -- a distance of 180 mi. Because he is flying into a head wind, the trip takes 2 hours. On the way back, the wind is blowing the same speed, so it takes an 1 hour and 48 mins. What is his speed in still air and how fast is the wind blowing?
Answer by Alan3354(69443)

A man flies an airplane from Fargo to Bismark -- a distance of 180 mi. Because he is flying into a head wind, the trip takes 2 hours. On the way back, the wind is blowing the same speed, so it takes an 1 hour and 48 mins. What is his speed in still air and how fast is the wind blowing?
---------------
Upwind, the groundspeed = 180/2 = 90 mph
Downwind, = 180/1.8 = 100 mph
The "speed in still air" is the airspeed, and it's the average = 95 mph
The windspeed is the difference = 5 mph